What is the slope of line A?
3x-7y=-3

Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

3x - 7y = -3
-7y = -3x - 3 {subtracted 3x from both sides}
y = (3/7)x + 3/7 {divided both sides by -7}

slope is 3/7
www.algebrahouse.com